volume 20 distilled 20 acetic HCL 20 20 NaOAc 20 naoh solution A 20 predict ph 7 4 .2 1 9 12 10.00 ml of 7 0.2 acetic acid and 10.00 ml 0.2 m sodium acetate
media%2F31b%2F31bf1264-74f5-42d8-9c4d-02
media%2Fff8%2Fff838a6d-d76d-4101-95c9-a2
Answer #1

Theoretical pH of the substances

substance M formula theoretical pH

acetic acid 0.2 pH = 1/2[pka -logC] 2.719

HCl 0.1 pH = -log[H+] 1.00   

NaOAc 0.2 pH = 1/2[pkw+pKa +logC] 9.02

NaOH 0.1 pH = 14 - pOH 13

solution A [acid] =0.1

[conj. base] =0.1 pH = pka + log [conj.base]/[acid] 4.74